OpenStudy (anonymous):

I have the answer can someone tell me what they got? Solve for x (3-x/3)+2=3x

OpenStudy (whpalmer4):

I've got an answer, let's see yours first...

OpenStudy (anonymous):

I got 2.6

OpenStudy (whpalmer4):

Rats. \[(3-\frac{x}3) + 2 = 3x\]Multiply everything by 3 \[9 - x + 6 = 9x\]\[x=\]?

OpenStudy (anonymous):

1.5

OpenStudy (whpalmer4):

there you go. And let's check it: 3-1.5/3 + 2 = 3(1.5) 2.5 + 2 = 4.5

OpenStudy (anonymous):

its right thanks

OpenStudy (whpalmer4):

check your answers like that and you won't have to ask us if you got the right answer...

OpenStudy (anonymous):

ok